The steps from this elog were followed.
In addition, I did a burt restore of c1sus, c1mcs.
I then swapped all the gain settings from ITMX to ITMY, and reenabled the watchdogs.
I did some basic kick tests (1000 counts into UL coil) and confirmed channels like C1:SUS-ITMX_ULPD_VAR (watchdogs mV readback) corresponded to the correct optic. I also checked that the POS, PIT, YAW, SIDE produced reasonable damping when engaged.